OpenStudy (anonymous):

A right triangle has a hypotenuse of 50 feet and a leg of 14 feet. Which is the length of the other leg?

OpenStudy (anonymous):

hypotenuse ^2=side1^2+side2^2 50^2=14^2+c^2 find c

OpenStudy (anonymous):

I don't know how

OpenStudy (anonymous):

c^2=50^2-14^2=2500-196=?.... can u do

OpenStudy (anonymous):

umm no.

OpenStudy (anonymous):

c^2=2304 c=r00t(2304)=48